# XL-Instruct: Synthetic Data for Cross-Lingual Open-Ended Generation

### v1

@@ 4 The XL-Instruct Pipeline
To address this gap, we introduce the XL-Instruct pipeline (illustrated in Figure 1), designed to create cross-lingual synthetic instructions from a given seed corpus. At this point, we highlight two important considerations. First, unlike related work (Li et al., 2024a), we select an English seed corpus for data generation. Our rationale here is to optimize the quality of synthetic data produced by teacher LLMs, as we hypothesize that these models should be more proficient at generating diverse and high-quality synthetic data in English. MT is employed only in the final stages, thereby minimizing noise propagation. Second, we exclusively utilize open-weight models with permissible licenses to generate synthetic data, aligning with our objective of releasing a fully permissible, open-source dataset.
We outline the four stages of the XL-Instruct pipeline in detail below:
@@
    Stage 1: Reverse Instructions Given a passage from our seed data, we ask a teacher LLM to generate an instruction for which this passage would be a valid response.
@@
    Stage 2: Refinement: Then, we ask the teacher to reword the question and response such that it follows four manually-defined criteria.
@@
    Stage 3: Response Translation Next, we translate the refined response, leveraging one or more translation LLMs, to the target language.
@@
    Stage 4: Filtering To ensure we use the highest quality targets, we use MT Quality Estimation (QE) models to select the best response translations in our dataset.
We conduct Supervised Fine-Tuning (SFT) on this synthetic dataset. We detail the minutiae of each stage below.
@@ 4.1 Stage 1: Question Generation
First, we sample an English passage from our seed corpus, CulturaX (Nguyen et al., 2024). Then, we ask a teacher LLM - here, Qwen 2.5 72B (Yang et al., 2024) - to produce an instruction for which the sampled sentence would be a valid response. Prompting in English allows us to leverage the teacher model directly without requiring the additional fine-tuning employed in previous work (Li et al., 2024a). To convert the LLM response to a cross-lingual instruction, we add a prompt to “Respond in {lang}” where {lang} is the target language of interest. This creates a potential synthetic instruction for SFT — the desired output for which would be the response translated to this language (see Section 4.3).
@@ 4.2 Stage 2: Refinement
    Question Self-Sufficiency: The question should be clear, unambiguous, and should not require any additional information/context to produce the given response.
@@
    Response Naturalness: The response should be ‘natural-sounding’ as an LLM output — in terms of fluency, neutrality and objectivity, and consistency with the tone and style of LLM-generated responses.
@@
    Response Precision: The response should be topically relevant, factually accurate and should directly answer the question. This can be thought of as analogous to precision since it tries to assess how much of the information contained in the response is actually relevant, necessary and true.
@@
    Response Informativeness: The response should be informative and helpful, and should contain enough justification and explanation to make it useful to an end user. This is similar to recall, as it evaluates how much of the relevant and useful information for the response is actually provided.
We provide all four criteria and their explanations in a prompt and ask the teacher to refine the (question, response) pair. We also instruct the model to ensure the reworded response is grounded in the original one, and request it to not add any of its own knowledge — since we want to avoid excessive teacher distillation and ensure our synthetic data is grounded in the seed corpus to optimize for diversity.
@@ 4.3 Stage 3: Response Translation
Next, the response must be machine translated into the target language. Since document-level MT by open LLMs is currently unreliable due to limited exploration, scarce datasets, and hallucination risks, we instead use sentence-level translation. We sentence-split using Segment Any Text (Frohmann et al., 2024) and generate translations in one of two ways:
@@
    Best-of-k: We obtain $k$ translations from $k$ different LLMs for each sentence, and choose the one with the best QE score as the final translation.
For QE, we use the WMT’23 CometKiwi-XL model (Rei et al., 2023), which obtained state-of-the-art (SOTA) scores in the WMT 2023 QE Shared Task (Blain et al., 2023). For MT, we use EuroLLM 9B Instruct (Martins et al., 2024) in the ‘naive’ case due to its strong MT capabilities, while in the ‘Best-of-k’ setting, we set $k=3$ and sample among EuroLLM 9B Instruct, Mistral Small 24B Instruct (Mistral AI Team, 2025) and Gemma2 27B Instruct (Team et al., 2024). Finally, we add a postprocessing step that involves creating the response translation by substituting each sentence-level translation in-place in the original response. This helps us retain formatting (like paragraph separators, bullet points etc.) that is critical to response quality.
@@ 4.4 Stage 4: Filtering
Finally, to ensure that we select high-quality targets during fine-tuning, we compute sentence-level QE scores using the WMT’23 CometKiwi-XL model, by comparing each source sentence in a given response and its translation. We average these QE scores across the entire passage to obtain the final passage-level score. Then, we sort all responses in descending order and filter the last 20% of dataset – creating a final dataset of about 32K instructions. We provide the prompts used in each stage at https://tinyurl.com/xli-prompts (to be uploaded to GitHub on acceptance)
@@ 5 Experiments on XL-AlpacaEval: Boosting Cross-Lingual Generation
We conduct SFT of two base (EuroLLM 9B, Qwen2.5 7B) and three instruction-tuned (EuroLLM 9B Instruct, Qwen2.5 7B Instruct and Aya Expanse 8B) models. We choose EuroLLM and Qwen since they relatively underperform on the XL-AlpacaEval benchmark (Table 1), leaving significant scope for improvement. We also experiment with Aya Expanse since it leads the benchmark, and we are interested in seeing how much further it could be improved. Unfortunately, Aya Expanse does not have a base model released, so we are unable to experiment with it.
@@
We fine-tune all models for 1 epoch using low-rank adaptation (LoRA, Hu et al., 2022) with rank 8 matrices applied to query and value projections. Training used a cosine learning rate scheduler with a peak learning rate of 1e-4 and 3% warmup steps. We employed bf16 mixed-precision training with batch size 8, and fix the random seed as 1 for reproducibility.
Main Results
@@
In Figure 2, we show the improved average win rates obtained by fine-tuning with various amounts of XL-Instruct data. For base LLMs, performance steadily improves with data scale until 8K training examples, at which point we observe saturation. Interestingly, EuroLLM consistently outperforms Qwen. The reason for this is revealed from the language-specific scores (Table 6) – the largest improvements are yielded for the languages the model is pre-trained on. For EuroLLM, which includes all 8 languages in its pre-training, there are significant, consistent, and relatively uniform gains across all languages, including the least-resourced ones. Qwen, which chiefly supports high-resource languages like Chinese, Portuguese and German, gains the most for these pairs but shows relatively smaller improvements for the others. This suggests that while post-training with XL-Instruct can yield stable improvements across multiple languages, multilingual pre-training is crucial for best performance.
@@
For instruction-tuned models, the saturation occurs much sooner—at around 2K instructions for EuroLLM-9B-Instruct and, only 32 instructions for the Qwen and Aya Expanse models! This can be explained by the fact that the latter two models have also undergone post-training with Preference Optimization—therefore, task-specific SFT at scale might lead to overfitting and deteriorated performance. EuroLLM has only been exposed to SFT data, and can therefore be trained for longer. In all cases, though, one observes consistent and significant improvements across all languages—with only 32 examples, we can push the win rates of Aya Expanse 8B to $\sim 65\%$ for its supported languages of Portuguese, German and Chinese (Table 6).
Finally, we show in the Appendix (Table 7) how XL-Instruct can also boost zero-shot cross-lingual performance, ie. even for languages not included during SFT.
Fine-Grained Evaluation
@@
Beyond win rates that depend solely on pairwise comparisons, we are also interested in evaluating how well the produced cross-lingual generations improve on an absolute scale, on human-desired criteria. To achieve this, we take inspiration from recent works that define customised, task-specific metrics and use LLM-as-a-judge for producing scores on a Likert scale – achieving significant correlations with human ratings on evaluation of tasks like summarization (Liu et al., 2023), retrieval (Upadhyay et al., 2024), story generation (Chiang & Lee, 2023), Machine Translation (Kocmi & Federmann, 2023b; a) and open-ended generation (Kim et al., 2023). In particular, Kim et al. (2023) showed that using clearly defined rubrics can result in Spearman correlations up to 0.87 with human preferences for open-ended generation. Inspired by this, we propose four criteria pertinent to the task of cross-lingual generation: Objectivity, Naturalness, Informativeness, and Precision. We define detailed rubrics for each metric and provide well-defined criteria for mapping output quality to scores on a scale of 1-5. We include these rubrics in the context of a prompt, and ask GPT-4o to score cross-lingual generations of EuroLLM 9B, EuroLLM 9B Instruct and EuroLLM 9B XL-Instruct (the best model from Figure 2(a), which is fine-tuned with LoRA on 8K examples). We provide detailed evaluation prompts and rubrics at https://tinyurl.com/xl-gen-eval (to be uploaded to GitHub on acceptance).
Figure 3: Evaluations on fine-grained quality metrics by GPT-4o on a scale of 1-5. For EuroLLM 9B XL-Instruct, we choose the best model from Figure 2(a). Despite being LoRA fine-tuned on only 8K instructions, it performs comparably or slightly better than EuroLLM 9B Instruct which has been fully fine-tuned on a 2M examples dataset. These scores are also tabulated in Table 8.
@@
We show our results in Figure 3, which provides the macro-averaged scores for each criterion per model. As one would expect, the untuned EuroLLM 9B base model achieves the worst scores on all metrics, with the EuroLLM-9B-Instruct model performing substantially better. We note that the XL-Instruct model performs comparably to or marginally better than the EuroLLM-9B-Instruct model. This result is particularly impressive given the XL-Instruct model was trained using LoRA fine-tuning on only 8K synthetic samples, whereas the EuroLLM-9B-Instruct model was fully fine-tuned on 2M human and synthetically-generated examples. These results clearly demonstrate the effectiveness and high quality of the XL-Instruct dataset.
@@ 6 Experiments on m-AlpacaEval: Exploring Zero-Shot Transfer
Table 3: Win Rates of LLMs and their XL-Instruct fine-tuned counterparts on m-AlpacaEval against GPT-4o-Mini, with GPT-4o as the judge. For each model, we choose the best cross-lingual performing baseline from Figure 2 and evaluate transfer on m-AlpacaEval. Consistent improvement across all models and pairs shows strong zero-shot transfer from cross-lingual tuning, for both multilingual and English-only generation. Best scores are bolded and cells are highlighted proportionate to performance gain.
Having seen task-specific improvements, we now seek to evaluate the performance of models fine-tuned with XL-Instruct on multilingual and English open-ended generation, since these are arguably the more common use cases of LLMs. For this purpose, we first construct the m-AlpacaEval benchmark by machine translating the AlpacaEval test set into our 8 languages of interest, following related efforts to create m-ArenaHard (Dang et al., 2024). We use GPT-4o for translation of the prompts. The evaluation setup is similar to XL-AlpacaEval, wherein GPT-4o-Mini is the reference model and GPT-4o is the judge.
We show the results in Table 3, using the base and instruct models from Figure 2 along with the best-performing cross-lingual baselines. We observe significant and consistent zero-shot transfer across all models and languages. For multilingual generation, the gains are strongest for the languages a model is pretrained on, similar to our observations for cross-lingual generation. This is particularly evident in the Qwen and Aya models. EuroLLM Instruct, on the other hand, achieves stable performances across all languages and relatively strongest win rates for the lower-resourced languages. Interestingly, one can also see consistent gains in English-only generation, despite there being no English responses on the target side! This suggests that all of these models, which are sufficiently proficient in English, can learn improved response structure and formatting from the cross-lingual responses. These results are quite encouraging, since they suggest cross-lingual fine-tuning need not come at the cost of standard ‘monolingual’ generation performance – in fact, it can result in further boosts.

### v2

@@ 4 The XL-Instruct Data Synthesis Pipeline
To address this gap, we introduce the XL-Instruct pipeline to create cross-lingual synthetic instructions from a given seed corpus, as illustrated in Figure˜1. We highlight two important considerations. First, unlike related work (Li et al., 2024a), we seed from English data instead of using the target language corpora directly. Given teacher LLMs are more proficient in English than in a low-resource language, we hypothesize that more high-quality, yet diverse, synthetic data could be generated in English. Machine translation is employed only in the final stages, thereby minimizing noise propagation. Second, we exclusively utilize open-weight models with permissible licenses to generate synthetic data, aligning with our objective of releasing a fully public open-source dataset.
The XL-Instruct pipeline contains four stages:
@@
    Stage 1 Reverse Instructions: Given a passage from our seed data, we ask a teacher LLM to generate an instruction for which this passage would be a valid response.
@@
    Stage 2 Refinement: Next, we ask the teacher to reword the question and response pairs to follow four manually defined criteria.
@@
    Stage 3 Response Translation: Then, we translate the refined response to the target language, using one or more translation LLMs.
@@
    Stage 4 Filtering: Finally, to ensure we use the highest quality targets, we use translation quality estimation (QE) models to filter the dataset for the best translations.
After the data is synthesized, we conduct supervised fine-tuning (SFT) on it with a range of models. We detail the minutiae in each subsection below.
@@ 4.1 Stage 1: Question Generation
First, we sample an English passage from our seed corpus, CulturaX (Nguyen et al., 2024). Then, we ask a teacher LLM (Qwen 2.5 72B (Yang et al., 2024)) to produce an instruction for which the sampled sentence would be a valid response. Prompting in English allows us to leverage the teacher model directly without requiring the additional fine-tuning employed previously (Li et al., 2024a). This stage thus yields a synthetic English instruction, paired with the English seed passage as a response.
@@ 4.2 Stage 2: Refinement
    Question Self-Sufficiency: The question should be clear and unambiguous, and should not require any additional information or context to produce the given response.
@@
    Response Naturalness: The response should be ‘natural-sounding’ as an LLM output — in terms of fluency, neutrality, objectivity, and consistency with the tone and style of LLM-generated responses.
@@
    Response Precision: The response should be topically relevant, factually accurate, and should directly answer the question. This can be thought of as analogous to precision since it tries to assess how much of the information contained in the response is relevant, necessary, and true.
@@
    Response Informativeness: The response should be informative and helpful, and must contain enough justification and explanation to make it useful to an end user. This is similar to recall, as it evaluates how much of the relevant and useful information for the response is actually provided.
We provide all four criteria and their definitions in a prompt and ask the teacher to refine the (question, response) pair. We also instruct the model to ensure the reworded response is grounded in the original one, and request it not to add any of its own knowledge—in order to avoid excessive teacher distillation and to ensure our targets are grounded in the seed data we use.
(a) Base LLMs
(b) Instruction-Tuned LLMs
@@ 4.3 Stage 3: Response Translation
Now, we direct our focus towards converting the English question-response pair to a cross-lingual En-X one. Creating the cross-lingual instruction itself is easy — we simply add a prompt to “Respond in {lang}” where {lang} is the target language of interest. To create the target, the English response must be machine-translated into the target language. Since document-level MT by open LLMs is currently unreliable due to limited exploration, scarce datasets, and hallucination risks, we use sentence-level translation instead. We sentence-split using Segment Any Text (Frohmann et al., 2024) and generate translations in one of two ways:
@@
    Best-of-k: We obtain $k$ translations from $k$ different LLMs for each sentence, and choose the one with the best QE score.
For QE, we use the WMT’23 CometKiwi-XL model (Rei et al., 2023), which obtained state-of-the-art scores in the WMT 2023 QE Shared Task (Blain et al., 2023). For translation, we use EuroLLM 9B Instruct (Martins et al., 2024) in the “naive” case due to its strong translation capabilities, while in the “best-of-k” setting, we set $k=3$ and sample among EuroLLM 9B Instruct, Mistral Small 24B Instruct (Mistral AI Team, 2025), and Gemma2 27B Instruct (Team et al., 2024). Finally, to create the translated response, we substitute each sentence in the original response with its sentence-level translation. This helps us retain formatting like paragraph separators, bullet points, etc, that is critical to response quality.
@@ 4.4 Stage 4: Filtering
Finally, to ensure that we select high-quality targets during fine-tuning, we compute sentence-level QE scores using the WMT’23 CometKiwi-XL model by comparing each source sentence in a given response and its translation. We average these QE scores across the entire passage to obtain the final passage-level score. Then, we sort all responses in descending order and filter the last 20% of the dataset – creating a final dataset of about 32K instructions. We release the prompts used in each stage on Hugging Face [^3] https://huggingface.co/datasets/viyer98/xl-instruct/resolve/main/data_prompts.py.
@@ 5 Experiments on XL-AlpacaEval: Boosting Cross-Lingual Generation
We conduct SFT of two base (EuroLLM 9B, Qwen2.5 7B) and three instruction-tuned (EuroLLM 9B Instruct, Qwen2.5 7B Instruct, and Aya Expanse 8B) models. We choose EuroLLM and Qwen since they relatively underperform on the XL-AlpacaEval benchmark (Table 1), leaving significant scope for improvement. We also experiment with Aya Expanse since it leads the benchmark, and we are interested in seeing how much further it could be improved. Unfortunately, Aya Expanse does not have a base model released, so we are unable to experiment with it.
@@
We fine-tune all models for 1 epoch using low-rank adaptation (LoRA, Hu et al., 2022) with rank 8 matrices applied to query and value projections. We also tune the input and output embeddings. Training used a cosine learning rate scheduler with a peak learning rate of 1e-4 and 3% warmup steps. We used bf16 mixed-precision training with batch size 8, and fixed the random seed at 1 for reproducibility. All experiments were run on 4 Nvidia GeForce RTX 3090 GPUs, each with 24 GB VRAM.
Model Avg fra fin tur
EuroLLM 9B 7.80 9.69 9.78 3.94
EuroLLM 9B Instruct 14.08 19.39 14.05 8.81
EuroLLM 9B XL-Instruct (best) 20.62 25.80 22.72 13.33
5.1 Main Results
@@
In Figure˜2, we report XL-AlpacaEval win rates on fine-tuning with various amounts of XL-Instruct data. We observe that for base LLMs, performance steadily improves with data scale. Qwen advances from a win rate of 5.8% to 13.89% against GPT-4o-mini, while EuroLLM achieves an even larger boost, going from 7.36% to as high as 21.89% on SFT with 8K instructions. We report language-specific scores in Table 9 and observe that while there are consistent gains for all languages, the largest gains are on an LLM’s pre-training language. Since EuroLLM includes all 8 XL-AlpacaEval languages in its pre-training, it observes large gains per language, leading to a much better overall average score. Qwen, which chiefly supports high-resource languages like Chinese, Portuguese, and German, gains the most for these pairs but shows relatively smaller improvements for others. This suggests that while post-training with XL-Instruct can yield stable improvements across multiple languages, multilingual pre-training is crucial for best performance.
@@
We also observe consistent improvements when fine-tuning instruction-tuned LLMs (Figure˜2(b)). Unlike base LLMs, the saturation occurs sooner here—at around 2K instructions for EuroLLM-9B-Instruct and, only 32 instructions for the Qwen and Aya Expanse models! This is likely because the latter two models have also undergone Preference Optimization, and task-specific SFT at scale might lead to overfitting and deteriorated performance. EuroLLM, on the other hand, has only undergone SFT, and can therefore be trained for longer.
Moreover, we report full results in Section˜B.1, where we observe consistent and major gains across all languages (Table 9). These results are particularly noteworthy given the modest training costs—low-rank fine-tuning with a few thousand instructions. Moreover, with only 32 examples, Aya Expanse achieves a win rate boost from ~57% to ~65% for its supported languages, Portuguese, German, Hindi, and Chinese (Table˜9). Lastly, we also show in Table˜3 in the Appendix how XL-Instruct can also boost zero-shot cross-lingual performance, i.e. even for languages not included in SFT.
Zero-Shot Results
Moreover, we show in Table 3 zero-shot cross-lingual performance after fine-tuning with XL-Instruct. We choose French (fra), Finnish (fin), and Turkish (hind), which are in EuroLLM’s pre-training. We see huge gains in win rates, largely outperforming even the official EuroLLM 9B Instruct. This shows that even if done only for a few languages, XL-Instruct can result in significant transfer that improves performance in others. We hypothesize that this is likely because the model is able to learn formatting, response structure, etc. from this process, which also supports the boosts in English generation one observes in Table 6.
5.2 Fine-Grained Evaluation
@@
Beyond win rates that focus solely on pairwise comparisons, we are also interested in evaluating how well the produced cross-lingual generations improve on an absolute scale, on human-desired criteria. To achieve this, we take inspiration from recent works that define customised, task-specific metrics and use LLM-as-a-Judge for producing scores on a Likert scale, which achieves strong correlations with human ratings on the evaluation of summarization (Liu et al., 2023), retrieval (Upadhyay et al., 2024), story generation (Chiang and Lee, 2023), translation (Kocmi and Federmann, 2023b, a), and open-ended generation (Kim et al., 2023). In particular, Kim et al. (2023) showed that using clearly defined rubrics can result in up to 0.87 Spearman correlations with human preferences for open-ended generation. Inspired by this, we propose four criteria pertinent to the task of cross-lingual generation: Objectivity, Naturalness, Informativeness, and Precision. We define detailed rubrics for each metric and provide well-defined criteria for mapping output quality to scores on a scale of 1-5. We include these rubrics in the context of a prompt, and ask GPT-4o to score cross-lingual generations of EuroLLM 9B, EuroLLM 9B Instruct, and EuroLLM 9B XL-Instruct (the best model from Figure 2(a), which is fine-tuned with LoRA on 8K examples). We provide detailed evaluation prompts and rubrics on Hugging Face [^4] https://huggingface.co/datasets/viyer98/xl-instruct/resolve/main/eval_prompts.py.
Model Avg Precision Informativeness Naturalness Objectivity
EuroLLM 9B 2.43 2.52 2.69 2.25 2.27
EuroLLM 9B Instruct (official) 3.56 3.68 3.80 3.54 3.23
EuroLLM 9B XL-Instruct (our best) 3.60 3.63 3.88 3.64 3.24
Table 4: Performance of EuroLLM 9B models evaluated by Precision, Informativeness, Naturalness, and Objectivity.
Model Avg fin hin tur
EuroLLM 9B + X-Instruction (full 1M) 9.73 14.35 7.22 7.63
EuroLLM 9B + X-Instruction (40K) 10.44 13.69 8.76 8.86
EuroLLM 9B + XL-Instruct (naive, 40K) 12.06 15.30 10.9 9.98
EuroLLM 9B + XL-Instruct (best of 3, 40K) 17.82 23.15 15.8 14.52
Table 5: Results from EuroLLM 9B fine-tuned on data from X-Instruction (Li et al., 2024a) and XL-Instruct (ours).
@@
We list rubric-based evaluation results in Table˜4, which provides the macro-averaged scores across criterion and model. As expected, the raw EuroLLM 9B base model achieves the worst scores on all metrics, with the EuroLLM-9B-Instruct model performing substantially better. We note that the XL-Instruct model performs comparably to or marginally better than EuroLLM-9B-Instruct. This result is particularly impressive given the XL-Instruct baseline was trained using LoRA fine-tuning on only 8K synthetic samples, whereas the EuroLLM-9B-Instruct was fully fine-tuned on a mix of 2M human and synthetic examples. These results clearly demonstrate the effectiveness and high quality of the XL-Instruct dataset.
5.3 Comparison to Previous Work
We also compare the efficacy of XL-Instruct with its most similar work—the only cross-lingual open-ended generation dataset we are aware of: X-Instruction (Li et al., 2024a). We base all comparisons on their public data, [^5] https://huggingface.co/datasets/James-WYang/X-Instruction using Hindi (hin), Finnish (fin), and Turkish (tur), because these are supported by EuroLLM and are also available in X-Instruction. We also generate XL-Instruct data in these languages, by redoing the XL-Instruct pipeline (Section 4) from Stage 3 (Response Translation) for these languages. We LoRA fine-tune EuroLLM 9B on various X-Instruction and XL-Instruct datasets. For the former, we use both the entire 1M-sized dataset available for these languages (in total) and a 40K instructions subset, which is more comparable to our XL-Instruct baselines. For XL-Instruct, we train two baselines—one trained on “naive” translations (i.e., using only EuroLLM 9B Instruct) and another using a “best-of-3” method (refer to Section˜4.3 for a detailed explanation).
We see that both XL-Instruct baselines significantly outperform X-Instruction, with our best model achieving a 70.68% improvement over the latter—showcasing the relative superiority of our pipeline. This also suggests it might be more effective to prompt a teacher model in English due to inherently superior capabilities, and we hypothesize it might allow for greater quality and diversity in responses, as well as allow for more complex operations like refinement following specifically defined, custom criteria.
@@ 6 Experiments on m-AlpacaEval: Exploring Zero-Shot Transfer
Table 6: Win Rates of LLMs and their XL-Instruct fine-tuned counterparts on m-AlpacaEval against GPT-4o-mini, judged by GPT-4o. For each model, we choose the best cross-lingual performing baseline from Figure 2 and evaluate transfer on m-AlpacaEval. Consistent improvement across all models and pairs shows strong zero-shot transfer from cross-lingual tuning, for both multilingual and English-only generation. Best scores are bolded and cells are highlighted proportionate to performance gain.
Having seen task-specific improvements, we now seek to evaluate the zero-shot performance of models fine-tuned with XL-Instruct on multilingual and English open-ended generation, since these are arguably the more common use cases of LLMs. For this purpose, we first construct the m-AlpacaEval benchmark by machine translating the AlpacaEval test set into our 8 languages of interest, following similar efforts to create m-ArenaHard (Dang et al., 2024). We use GPT-4o for translation of the prompts. The evaluation setup is similar to XL-AlpacaEval, wherein GPT-4o-mini is used as the reference model and GPT-4o is used as the judge.
We present our results in Table 6, for the base and instruct versions of the LLMs from Figure 2, alongside their best-performing XL-Instruct-tuned counterparts. We observe significant and consistent zero-shot transfer across all models and languages. For multilingual generation, the gains are strongest for the languages a model is pre-trained on, similar to our observations for cross-lingual generation. This is particularly evident in the Qwen and Aya models. EuroLLM Instruct, on the other hand, achieves stable performances across all languages and relatively strongest win rates for the lower-resourced languages. Interestingly, we also note consistent gains in English-only generation, despite there being no English responses on the target side! This suggests that all of these models, trained heavily on English, can learn preferred response structure and formatting from cross-lingual tuning. These results are quite encouraging, since they suggest cross-lingual fine-tuning need not come at the cost of standard “monolingual” generation performance—on the contrary, it can result in further boosts.